ISOMERS AND MICROSOLVATION IN SiOH^{+}-Ar_{n} COMPLEXES (n=1-10)

\underline{OTTO DOPFER}, ROUSLAN V. OLKHOV, and SERGEY A. NIZKORODOV, Institute for Physical Chemistry, University of Basle, Klingelbergstrasse 80, CH-4056 Basel, Switzerland.

Infrared photodissociation spectra of mass selected SiOH^{+}-Ar_{n} ionic complexes have been recorded in the vicinity of the OH stretch vibration in a tandem mass spectrometer. Two isomers are identified in the spectrum of the dimer (n=1) via rotational and vibrational analysis and comparison with ab initio calculations: a linear proton bound dimer and a T-shaped complex. Though the spectra of larger clusters (n=2-10) display only vibrational resolution, the analysis of the systematic complexation induced frequency shifts provides a detailed picture of the cluster growth, including the formation of solvation rings and the existence of various isomers^{a}.

aR. V. Olkhov, S. A. Nizkorodov and O. Dopfer, Chem. Phys. \underline{239}, 393, 1998.